OpenStudy (anonymous):

Marcus is selling t-shirts at the State Fair. He brings 200 shirts to sell. He has long sleeve and short sleeved T-shirts for sale. On the first day of the fair, he sells 1/2 of his long sleeved T-shirts and 1/3 of his short sleeved t-shirts for a total of 80 t-shirts sold. How much of each type of shirt did Marcus bring to the fair ?

OpenStudy (kropot72):

Let x = number of long sleeve shirts Let y = number of short sleeve shirts x + y = 200 x/2 + y/3 = 80 Now solve this pair of simultaneous equations to find x and y.

OpenStudy (anonymous):

how would I solve it ?

OpenStudy (kropot72):

x + y = 200. Therefore x = 200 - y ...................(1) x/2 + y/3 = 80 .............(2) Substitute the value for x in (1) into equation (2) and solve for y.